Find the volume of a frustum of a regular square pyramid if the base edges are 7 cm and 19 cm, and one lateral edge is inclined at an angle of 60 degrees with the lower base.

 


1
Expert's answer
2021-02-24T13:49:50-0500

Volume of a frustrum = Volume of the larger pyramid - Volume of the smaller pyramid

V = 1/3(BASE AREA*HEIGHT) - 1/3(base area*height)

Please follow the attached image for guidance

sin 450 = x1cm/19

x1cm = 19*sin 450 = 13.44 cm

Now calculating the H, Height of the larger pyramid we have;

Tan 600 = H/13.44

H = 13.44 *Tan 600 = 23.28 cm


sin 450 = x2cm/7

x2cm = 7*sin 450 = 4.95 cm

Now calculating the h, Height of the larger pyramid we have;

Tan 600 = h/4.95

h = 4.95 *Tan 600 = 8.57 cm

Therefore the volume of the frustrum is calculated as below;

(1/3*192*23.28) - (1/3*72*8.57)

2801.36 - 139.98

V = 2661.38 cm3